1. Find a 3 × 3 matrix A that has eigenvalues λ = 0, 6, − 6 with corresponding eigenvectors

0 0 0
1 , -1 , 1
-1 1 1

2. Let Upper T ⁢ colon Upper M Subscript 22 Baseline ⁢ right-arrow Upper M Subscript 22 be the dilation operator with factor k = 3.

(a) Find Upper T left-parenthesis Start 2 By 2 Matrix 1st Row 1st Column 1, 2nd Column 4 2nd Row 1st Column -7, 2nd Column 4 End Matrix right-parenthesis = ????

(b) Find the rank and nullity of T .

3. Let Upper T ⁢ colon Upper P Subscript 2 Baseline ⁢ right-arrow Upper P Subscript 2 be the contraction operator with factor k ⁢ equals StartFraction 1 Over 6 EndFraction Number .

(a) Find Upper T left-parenthesis 1 ⁢ plus 6 x ⁢ plus 12 x Superscript 2 Baseline right-parenthesis Number .

(b) Find the rank and nullity of T.
1
Expert's answer
2020-01-29T13:24:14-0500

A diagonalizable matrix is diagonalized by a matrix of its eigenvectors. So "A=P\Lambda P^{-1 }" where "P" is the matrix whose columns are the eigenvectors of "A" and "\Lambda" is a diagonal matrix whose diagonal entries are the eigenvalues of "A."

"P=\begin{pmatrix}
 1 & 0 & 0 \\
 1 & -1 & 1 \\
 -1 & 1 & 1
\end{pmatrix} , \Lambda=\begin{pmatrix}
 0 & 0 & 0 \\
 0 & 6 & 0 \\
 0 & 0 & -6
\end{pmatrix}"

Find "P^{-1}"


"\begin{vmatrix}
 1 & 0 & 0 \\
 1 & -1 & 1 \\
 -1 & 1 & 1
\end{vmatrix}=1\begin{vmatrix}
 -1 & 1 \\
 1 & 1
\end{vmatrix}=1(-1-1)=-2\not=0"

"C_{11}=(-1)^{1+1}\begin{vmatrix}
 -1 & 1 \\
 1 & 1
\end{vmatrix}=-2,"

"C_{12}=(-1)^{1+2}\begin{vmatrix}
 1 & 1 \\
 -1 & 1
\end{vmatrix}=-2,"

"C_{13}=(-1)^{1+3}\begin{vmatrix}
 1 & -1 \\
 -1 & 1
\end{vmatrix}=0,"

"C_{21}=(-1)^{2+1}\begin{vmatrix}
 0 & 0 \\
 1 & 1
\end{vmatrix}=0,"

"C_{22}=(-1)^{2+2}\begin{vmatrix}
 1 & 0 \\
 -1 & 1
\end{vmatrix}=1,"

"C_{23}=(-1)^{2+3}\begin{vmatrix}
 1 & 0 \\
 -1 & 1
\end{vmatrix}=-1,"

"C_{31}=(-1)^{3+1}\begin{vmatrix}
 0 & 0 \\
 -1 & 1
\end{vmatrix}=0,"

"C_{32}=(-1)^{3+2}\begin{vmatrix}
 1 & 0 \\
 1 & 1
\end{vmatrix}=-1,"

"C_{33}=(-1)^{3+3}\begin{vmatrix}
 1 & 0 \\
 1 & -1
\end{vmatrix}=-1."


"C=\begin{pmatrix}
 -2 & -2 & 0 \\
 0 & 1 & -1 \\
 0 & -1 & -1
\end{pmatrix}"

The adjount matrix is:


"\begin{pmatrix}
 -2 & 0& 0 \\
 -2 & 1 & -1 \\
 0 & -1 & -1
\end{pmatrix}""P^{-1}=\begin{pmatrix}
 1 & 0& 0 \\
 1 & -0.5 & 0.5 \\
 0 & 0.5 & 0.5
\end{pmatrix}"

"P\Lambda=\begin{pmatrix}
 1 & 0 & 0 \\
 1 & -1 & 1 \\
 -1 & 1 & 1
\end{pmatrix}\begin{pmatrix}
 0 & 0 & 0 \\
 0 & 6 & 0 \\
 0 & 0 & -6
\end{pmatrix}="

"=\begin{pmatrix}
 0 & 0 & 0 \\
 0 & -6 & -6 \\
 0 & 6 & -6
\end{pmatrix}"

"A=P\Lambda P^{-1 }=\begin{pmatrix}
 0 & 0 & 0 \\
 0 & -6 & -6 \\
 0 & 6 & -6
\end{pmatrix} \begin{pmatrix}
 1 & 0& 0 \\
 1 & -0.5 & 0.5 \\
 0 & 0.5 & 0.5
\end{pmatrix}="

"=\begin{pmatrix}
 0 & 0 & 0 \\
 -6 & 0 & -6 \\
 6 & -6 & 0
\end{pmatrix}"


"A=\begin{pmatrix}
 0 & 0 & 0 \\
 -6 & 0 & -6 \\
 6 & -6 & 0
\end{pmatrix}"


2.

(a) Let "T:M_{22}\to M_{22}" be the dilation operator with factor k = 3.

This means that "T(A)=3A \ \ \ \forall A\in M_{22}"


"A=\begin{pmatrix}
 1 & 4 \\
 -7 & 2
\end{pmatrix}"

Then


"T(A)=3A=3\begin{pmatrix}
 1 & 4 \\
 -7 & 2
\end{pmatrix}=\begin{pmatrix}
 3 & 12 \\
 -21 & 6
\end{pmatrix}"

"T(A)=\begin{pmatrix}
 3 & 12 \\
 -21 & 6
\end{pmatrix}"

(b) Find the rank and nullity of T.

As we know, nullity represents dimension of Kernel.

We have :


"B\in Ker{\{T\}}=>T(B)=0""=>3B=0=>Ker{\{T\}}=0"

Which means that


"nullity(T)=0"

"dim(M_{22})=rank(T)+nullity(T)""dim(M_{22})=4, nullity(T)=0"

Hence


"rank(T)=dim(M_{22})-nullity(T)=4-0=4""rank(T)=4"

3.

Let "T:P_{2}\to P_{2}" be the contraction operator with factor k = 1/6.

Find "T(1+6x+12x^2)"


"T(p(x))={1 \over 6}p(x),\ \forall p(x)\in P_2"

That gives us:


"T(1+6x+12x^2)={1 \over 6}(1+6x+12x^2)={1 \over 6}+x+2x^2""T(1+6x+12x^2)={1 \over 6}+x+2x^2"

(b) Find the rank and nullity of T.

As we know, nullity represents dimension of Kernel.

We have :


"p\in Ker{\{T\}}=>T(p(x))=0, \forall x""=>{1 \over 6}p(x)=0=>Ker{\{T\}}=0"

Which means that


"nullity(T)=0"

"dim(P_{2})=rank(T)+nullity(T)""dim(P_{2})=3, nullity(T)=0"

Hence


"rank(T)=dim(P_{2})-nullity(T)=3-0=3""rank(T)=3"